Identify the discontinuities , f(x)=(x^2-2x-3)/(x+1)
** please answer by tommorow!!
1
Expert's answer
2019-08-01T13:21:16-0400
"f(x)=\frac{x^2-2x-3}{x+1}=\frac{(x-3)(x+1)}{x+1}=x-3"

thus, this function is just a simple line and has a discontinuity at x = -1.